# Investigating Feature and Model Importance in Android Malware Detection: An Implemented Survey and Experimental Comparison of ML-Based Methods

### v3

@@ 3.6. Opcodes
3.7. Comparison of static modelling approaches
To identify the best combination of features, ML models and feature selection algorithms for static analysis, we carried out a Kruskal-Wallis test followed by a pairwise comparison between the modelling approaches from this section. The results indicate that five of the models have statistically better accuracies than the others. Three of these are API call models: SVMs trained using 10,000 features selected using Chi-square, RF models trained using 10,000 features selected using MI, and RF models trained using 6,443 features selected using variance threshold (Table 12). The others are the RF usage 5-opcode model, trained using 46,781 features selected using mutual information (Table 17) and DNN trained using API frequency (Table 7). Neither of these five modelling approaches appears to be significantly better than the others. This suggests that the best classes of features to use for static analysis are opcodes and API calls, and the best ML model to use is likely to be RF. However, the best choice of feature selection algorithm depends on the feature type and model type.
4. Dynamic Analysis
@@ 4.1. System Calls
Ananya et al. (Ananya et al., 2020) used sequences of system calls represented as unigrams, bigrams, and trigrams as the features of their models. They also used feature selection to select relevant features, and trained four ML models: linear regression, decision tree, RF and XGBoost. They used a dataset of 2,475 benign applications and 2,474 malware. The authors proposed two novel approaches to feature selection. The first, called SAILS, builds on the conventional feature selection algorithms mutual information, Chi-square and DFS. These are used to score the benign and malware features. SAILS then sorts the resulting scores in descending order, to create two sorted lists of benign and malware features. It then creates a final list that is the union of these two lists. The second feature selection algorithm proposed was WFS, which assigns weights to system calls. This is done by calculating the ratio of the number of occurrences of the system call in malware to the number of occurrences in the sum of benign and malware. In all the experiments, the authors used SAILS, WFS, mutual information, Chi-square and DFS to select relevant features. In our reimplementation, we also add PCC.
@@
| Unigram | ||||||||||||
| Classifier | Feature Selection | Accuracy | Precision | F1-Score | TPR | TNR | Feature Selection | Accuracy | Precision | F1-Score | TPR | TNR |
| SVM | - | - | - | - | - | - | Mutual information | 0.928 | 0.926 | 0.924 | 0.922 | 0.929 |
| Linear Regression | Mutual information | 0.977 | - | - | - | - | Mutual information | 0.921 | 0.922 | 0.919 | 0.915 | 0.923 |
| Random Forest | WFS | 0.972 | - | - | - | - | Mutual information | 0.934 | 0.935 | 0.933 | 0.931 | 0.936 |
| Decision Tree | Mutual information | 0.965 | - | - | - | - | Mutual information | 0.925 | 0.926 | 0.921 | 0.922 | 0.926 |
| XGBoost | WFS | 0.966 | - | - | - | - | Mutual information | 0.933 | 0.931 | 0.930 | 0.931 | 0.932 |
| kNN | - | - | - | - | - | - | Mutual information | 0.922 | 0.918 | 0.921 | 0.917 | 0.923 |
| Bigram | ||||||||||||
| Classifier | Feature Selection | Accuracy | Precision | F1-Score | TPR | TNR | Feature Selection | Accuracy | Precision | F1-Score | TPR | TNR |
| SVM | - | - | - | - | - | - | Mutual information | 0.941 | 0.939 | 0.942 | 0.938 | 0.943 |
| Linear Regression | Chi-square | 0.991 | - | - | - | - | Mutual information | 0.936 | 0.930 | 0.934 | 0.938 | 0.931 |
| Random Forest | Chi-square | 0.988 | - | - | - | - | Chi-square | 0.945 | 0.941 | 0.943 | 0.945 | 0.941 |
| Decision Tree | Chi-square | 0.977 | - | - | - | - | Mutual information | 0.931 | 0.929 | 0.929 | 0.931 | 0.929 |
| XGBoost | DFS | 0.994 | - | - | - | - | Chi-square | 0.943 | 0.943 | 0.942 | 0.939 | 0.944 |
| kNN | - | - | - | - | - | - | Chi-square | 0.932 | 0.929 | 0.931 | 0.926 | 0.931 |
| Trigram | ||||||||||||
| Classifier | Feature Selection | Accuracy | Precision | F1-Score | TPR | TNR | Feature Selection | Accuracy | Precision | F1-Score | TPR | TNR |
| SVM | - | - | - | - | - | - | Chi-square | 0.947 | 0.945 | 0.946 | 0.947 | 0.951 |
| Linear Regression | Mutual information | 0.994 | - | - | - | - | Mutual information | 0.932 | 0.932 | 0.934 | 0.937 | 0.932 |
| Random Forest | Mutual information | 0.982 | - | - | - | - | Chi-square | 0.951 | 0.951 | 0.950 | 0.949 | 0.952 |
| Decision Tree | WFS | 0.987 | - | - | - | - | Mutual information | 0.945 | 0.943 | 0.942 | 0.943 | 0.943 |
| XGBoost | Mutual information | 0.992 | - | - | - | - | Chi-square | 0.950 | 0.948 | 0.947 | 0.947 | 0.948 |
| kNN | - | - | - | - | - | - | Chi-square | 0.942 | 0.941 | 0.943 | 0.938 | 0.941 |
@@
We reimplemented one more study that built models solely from system call features, carried out by Malik et al. (Malik et al., 2019). This study is notable for using a different group of ML models and a different representation of the features. In particular, they used a 3-layer LSTM to classify malware based on the complete sequence of system calls used by an application, comparing this against a kNN model trained on a system call usage feature set. Both and the number of neurons per LSTM layer were optimised. Table 21 shows the results. Like Malik et al., we found that there was no benefit to using a more complex, slower to train, sequence-based model, since the best kNN model (with ) substantially out-performed the best LSTM (with 128 neurons per layer). However, our LSTM model performed much better than the model reported by Malik et al.
@@
4.2. API Calls
Perhaps an obvious question at this point is whether API calls, which we found to be the most effective class of features in models resulting from static analysis, would also be beneficial for dynamic analysis. Afonso et al. (Afonso et al., 2015) addressed this question to a certain extent by using information about both system calls and API calls in their models. Specifically, they used call frequencies as features, and a dataset of 2,968 benign and 4,552 malicious applications. We reimplemented their approach using SVM, RF, decision trees, kNN and Naïve Bayes. Afonso et al. also considered several other models, but those performed poorly in their study, so we did not reimplement them.
| Afonso et al. Results | Our Results | ||||||||||
| Random Forest | 0.968 | 0.975 | 0.968 | 0.961 | 0.976 | Random Forest | 0.962 | 0.968 | 0.964 | 0.959 | 0.969 |
| SVM | 0.951 | 0.953 | 0.952 | 0.949 | 0.959 | ||||||
| kNN | 0.946 | 0.943 | 0.946 | 0.943 | 0.948 | ||||||
| DT | 0.936 | 0.938 | 0.935 | 0.933 | 0.938 | ||||||
| Naïve Bayes | 0.857 | 0.860 | 0.856 | 0.851 | 0.862 | ||||||
@@ 4.3. Network Traffic
| TCP | DT | 0.982 | - | - | - | - | Random Forest | 0.973 | 0.967 | 0.977 | 0.973 | 0.969 |
| HTTP | DT | 0.997 | - | - | - | - | Random Forest | 0.963 | 0.961 | 0.965 | 0.965 | 0.961 |
| Network Feature | Classifier | Accuracy | Precision | F1-Score | TPR | TNR |
| HTTP | SVM | 0.941 | 0.942 | 0.941 | 0.943 | 0.946 |
| HTTP | Random Forest | 0.963 | 0.961 | 0.965 | 0.965 | 0.961 |
| HTTP | DT | 0.932 | 0.935 | 0.937 | 0.933 | 0.932 |
| HTTP | kNN | 0.951 | 0.956 | 0.955 | 0.949 | 0.959 |
| TCP + HTTP | SVM | 0.938 | 0.933 | 0.934 | 0.933 | 0.932 |
| TCP + HTTP | Random Forest | 0.952 | 0.953 | 0.956 | 0.955 | 0.951 |
| TCP + HTTP | DT | 0.929 | 0.926 | 0.925 | 0.922 | 0.925 |
| TCP + HTTP | kNN | 0.967 | 0.964 | 0.967 | 0.969 | 0.964 |
@@
Zulkifli et al. (Zulkifli et al., 2018) used DT with TCP-based features in their study. They used a small dataset of 500 benign and 200 malicious applications, the latter assembled from two existing malware collections (Drebin and contagiodumpset). Notably, all the applications were executed on a real Android device. They extracted the following features:
@@
To further investigate network traffic features, we reimplemented another study by Wang et al. (Wang et al., 2016) involving network data. Wang et al. extracted both TCP and HTTP features from network traffic. Wang et al. used the same set of TCP features as Zulkifli et al., with the exception of average packet size. The features they extracted from HTTP headers were:
@@
Table 26 shows the results, and compares them against those from the original study. For our reimplementation, Table 26 shows the results of the best ML models, which were RFs for both the HTTP and TCP feature sets. We observed no benefit to combining the two feature sets, but, for completeness, we also show results for the other models for both the HTTP and combined feature sets in Table 27. Whilst the results show good discrimination between benign applications and malware, we did not see the exceptionally high level of accuracy reported by Wang et al. for TCP features, and in fact found these to be less discriminative than HTTP features. These conclusions are supported by a positive Kruskal-Wallis test (H = 14.29, p 0.05) and Dunn’s tests.
@@
4.4. Comparison of dynamic modelling approaches
@@
5. Hybrid Analysis
| Kandukuru and Sharma Results | Our Results | ||||||||||
| DT | 0.956 | - | - | 0.95 | 0.96 | Random Forest | 0.943 | 0.941 | 0.945 | 0.948 | 0.940 |
| SVM | 0.933 | 0.936 | 0.935 | 0.939 | 0.937 | ||||||
| DT | 0.921 | 0.923 | 0.923 | 0.926 | 0.925 | ||||||
| kNN | 0.936 | 0.938 | 0.939 | 0.941 | 0.937 | ||||||
@@ 5. Hybrid Analysis
| Decision Trees | 0.919 | 0.919 | 0.917 | 0.923 | 0.917 |
| Classifier | No. of Features | Accuracy | Precision | F1-Score | TPR | TNR |
@@
We begin by reimplementing a study by Kandukuru and Sharma (Kandukuru and Sharma, 2017) that combined dynamic network traffic features with static permission features. The network traffic features included all seven of the TCP features listed in Section 4.3 and the authors used a malware dataset of 1,200 applications collected from 2010 to 2011 to train a decision tree model. In our reimplementation, we trained all four core ML models, using a dataset of feature vectors comprising 166 permissions and the seven TCP features. Table 28 compares their performance with Kandukuru and Sharma’s results. Since their performance is worse than models trained only on network features (see Tables tables 25 and 26), this suggests that there is no benefit to including information about static permissions.
To further investigate the use of hybrid features, we reimplemented a study based around two frequently used groups of static and dynamic features, permissions and system calls. Kapratwar et al. (Kapratwar et al., 2017) carried out experiments using these features both separately and in combination. In the case of system calls, a frequency-based representation was used. For each experiment, the authors trained RF, decision trees, Naïve Bayes, logistic regression, kNN and SVM. However, as with most past hybrid works, Kapratwar et al. used a small dataset of just 103 malware and 97 benign applications. Therefore, a reimplementation of this approach using a larger contemporary dataset is essential to understand the effect of combining two of the most prevalent features in Android malware detection. The dataset used for this reimplementation included 166 permissions and 599 system calls.
@@
6. Best-Performing Models
Table 31 lists the best classification models seen during this study. Two modelling approaches stand out in terms of the metric scores achieved when discriminating malware from benign applications. TCP network traffic models using random forests () achieved the highest scores in terms of accuracy, precision and TNR. Five-gram opcode models, again using random forests (), achieved the highest scores in terms of F1 and TPR. As noted in Section 2.3, TPR is particularly important for malware detection, since a benign application labelled as malware is less likely to cause as many security problems as malware labelled as benign. It is also notable that 11 of the 12 classifiers listed in Table 31 use random forests, suggesting that this is a good default choice within malware detection. This reflects the wider success of tree-based ensemble models, especially when working with pre-extracted features (Grinsztajn et al., 2022).
@@ 6. Best-Performing Models
A notable omission within this table is the presence of any deep learning models. In the majority of Android malware detection studies, features are pre-extracted. In this situation, there is often little benefit to using a deep learning model (Grinsztajn et al., 2022), since their strength lies in their ability to extract features from raw data. Deep learning models also bring risks of overfitting, responding to spurious patterns within data, and limited transparency (Lones, 2021), not to mention higher training and inference times — so in general are not likely to be a productive approach for malware detection unless they lead to a significant improvement in performance.
| ID | Feature Set | Type | Model | Feature Selection | Accuracy | F1-Score | Precision | TPR | TNR |
| API | Static | Random Forest | Mutual information | 0.967 | 0.969 | 0.967 | 0.963 | 0.970 | |
| API | Static | Random Forest | Variance threshold | 0.968 | 0.971 | 0.968 | 0.965 | 0.971 | |
| API | Static | Random Forest | Chi-square | 0.961 | 0.915 | 0.954 | 0.967 | 0.953 | |
| Permissions | Static | Random Forest | Mutual information | 0.930 | 0.932 | 0.929 | 0.931 | 0.931 | |
| Permissions | Static | Random Forest | Variance threshold | 0.949 | 0.948 | 0.950 | 0.951 | 0.949 | |
| Opcodes (5-gram) | Static | Random Forest | Mutual information | 0.973 | 0.979 | 0.975 | 0.965 | 0.979 | |
| System Calls (3-gram) | Dynamic | Random Forest | Chi-square | 0.951 | 0.951 | 0.950 | 0.949 | 0.952 | |
| System Calls + API Calls (Usage) | Dynamic | Random Forest | None | 0.961 | 0.964 | 0.963 | 0.957 | 0.965 | |
| Network Traffic (TCP) | Dynamic | Random Forest | None | 0.977 | 0.969 | 0.981 | 0.978 | 0.971 | |
| Network Traffic (TCP + HTTP) | Dynamic | kNN | None | 0.967 | 0.964 | 0.967 | 0.969 | 0.964 | |
| Network Traffic (TCP) + Permissions | Hybrid | Random Forest | None | 0.943 | 0.941 | 0.945 | 0.948 | 0.940 | |
| Permissions + System Calls | Hybrid | Random Forest | Mutual information | 0.948 | 0.948 | 0.949 | 0.935 | 0.961 |
7. Ensemble Models

### v4

@@ 5.6 Use and representation of opcode features
5.7 Comparison of static modelling approaches
To identify the best combination of features, ML models and feature selection algorithms for static analysis, we carried out a Kruskal-Wallis test followed by a pairwise comparison between the modelling approaches from this section. The results indicate that five of the models have statistically better accuracies than the others. Three of these are API call models: SVMs trained using 10,000 features selected using chi-square, RF models trained using 10,000 features selected using MI, and RF models trained using 6,443 features selected using variance threshold (Table 12). The others are the RF usage 5-opcode model, trained using 46,781 features selected using mutual information (Table 17) and DNN trained using API frequency (Table 7). Neither of these five modelling approaches appears to be significantly better than the others. Within the scope of our evaluation, this suggests that the best classes of features to use for static analysis are opcodes and API calls, and the best ML model to use is RF. However, the best choice of feature selection algorithm depends on the feature type and model type.
6 Reimplementation and extension of dynamic modelling approaches
@@ 6.1 Use of system call features
Ananya et al. [Ananya2020] used sequences of system calls represented as unigrams, bigrams, and trigrams as the features of their models. They also used feature selection to select relevant features, and trained four ML models: linear regression, decision tree, RF and XGBoost. They used a dataset of 2,475 benign applications and 2,474 malware. The authors proposed two novel approaches to feature selection. The first, called SAILS, builds on the conventional feature selection algorithms mutual information, chi-square and DFS. These are used to score the benign and malware features. SAILS then sorts the resulting scores in descending order, to create two sorted lists of benign and malware features. It then creates a final list that is the union of these two lists. The second feature selection algorithm proposed was WFS, which assigns weights to system calls. This is done by calculating the ratio of the number of occurrences of the system call in malware to the number of occurrences in the sum of benign and malware. In all the experiments, the authors used SAILS, WFS, mutual information, chi-square and DFS to select relevant features. In our reimplementation, we also add PCC.
@@
| Classifier | Selector | Accuracy | Precision | F1-Score | TPR | TNR | Selector | Accuracy | Precision | F1-Score | TPR | TNR |
| Features: Unigrams | ||||||||||||
| SVM | - | - | - | - | - | - | MI | 0.928 | 0.926 | 0.924 | 0.922 | 0.929 |
| Linear Regression | MI | 0.977 | - | - | - | - | MI | 0.921 | 0.922 | 0.919 | 0.915 | 0.923 |
| RF | WFS | 0.972 | - | - | - | - | MI | 0.934 | 0.935 | 0.933 | 0.931 | 0.936 |
| DT | MI | 0.965 | - | - | - | - | MI | 0.925 | 0.926 | 0.921 | 0.922 | 0.926 |
| XGBoost | WFS | 0.966 | - | - | - | - | MI | 0.933 | 0.931 | 0.930 | 0.931 | 0.932 |
| kNN | - | - | - | - | - | - | MI | 0.922 | 0.918 | 0.921 | 0.917 | 0.923 |
| Features: Bigrams | ||||||||||||
| SVM | - | - | - | - | - | - | MI | 0.941 | 0.939 | 0.942 | 0.938 | 0.943 |
| Linear Regression | CS | 0.991 | - | - | - | - | MI | 0.936 | 0.930 | 0.934 | 0.938 | 0.931 |
| RF | CS | 0.988 | - | - | - | - | CS | 0.945 | 0.941 | 0.943 | 0.945 | 0.941 |
| DT | CS | 0.977 | - | - | - | - | MI | 0.931 | 0.929 | 0.929 | 0.931 | 0.929 |
| XGBoost | DFS | 0.994 | - | - | - | - | CS | 0.943 | 0.943 | 0.942 | 0.939 | 0.944 |
| kNN | - | - | - | - | - | - | CS | 0.932 | 0.929 | 0.931 | 0.926 | 0.931 |
| Features: Trigrams | ||||||||||||
| SVM | - | - | - | - | - | - | CS | 0.947 | 0.945 | 0.946 | 0.947 | 0.951 |
| Linear Regression | MI | 0.994 | - | - | - | - | MI | 0.932 | 0.932 | 0.934 | 0.937 | 0.932 |
| RF | MI | 0.982 | - | - | - | - | CS | 0.951 | 0.951 | 0.950 | 0.949 | 0.952 |
| DT | WFS | 0.987 | - | - | - | - | MI | 0.945 | 0.943 | 0.942 | 0.943 | 0.943 |
| XGBoost | MI | 0.992 | - | - | - | - | CS | 0.950 | 0.948 | 0.947 | 0.947 | 0.948 |
| kNN | - | - | - | - | - | - | CS | 0.942 | 0.941 | 0.943 | 0.938 | 0.941 |
@@
We reimplemented one more study that built models solely from system call features, carried out by Malik et al. [Malik2019]. This study is notable for using a different group of ML models and a different representation of the features. In particular, they used a 3-layer LSTM to classify malware based on the complete sequence of system calls used by an application, comparing this against a kNN model trained on a system call usage feature set. Both and the number of neurons per LSTM layer were optimised. Table 21 shows the results. Like Malik et al., we found that there was no benefit to using a more complex, slower to train, sequence-based model, since the best kNN model (with ) substantially out-performed the best LSTM (with 128 neurons per layer). However, our LSTM model performed much better than the model reported by Malik et al.
@@
6.2 Use of dynamic API call features
Perhaps an obvious question at this point is whether API calls, which we found to be the most effective class of features in models resulting from static analysis, would also be beneficial for dynamic analysis. Afonso et al. [Afonso2015] addressed this question to a certain extent by using information about both system calls and API calls in their models. Specifically, they used call frequencies as features, and a dataset of 2,968 benign and 4,552 malicious applications. We reimplemented their approach using SVM, RF, decision trees, kNN and Naïve Bayes. Afonso et al. also considered several other models, but those performed poorly in their study, so we did not reimplement them.
| Afonso et al. Results | Our Results | |||||||||
| RF | 0.968 | 0.975 | 0.968 | 0.961 | 0.976 | 0.962 | 0.968 | 0.964 | 0.959 | 0.969 |
| SVM | - | - | - | - | - | 0.951 | 0.953 | 0.952 | 0.949 | 0.959 |
| kNN | - | - | - | - | - | 0.946 | 0.943 | 0.946 | 0.943 | 0.948 |
| DT | - | - | - | - | - | 0.936 | 0.938 | 0.935 | 0.933 | 0.938 |
| Naïve Bayes | - | - | - | - | - | 0.857 | 0.860 | 0.856 | 0.851 | 0.862 |
@@ 6.3 Use of network traffic features
| TCP | DT | 0.982 | - | - | - | - | RF | 0.973 | 0.967 | 0.977 | 0.973 | 0.969 |
| HTTP | DT | 0.997 | - | - | - | - | RF | 0.963 | 0.961 | 0.965 | 0.965 | 0.961 |
| Classifier | Accuracy | Precision | F1-Score | TPR | TNR |
| Features: HTTP | |||||
| SVM | 0.941 | 0.942 | 0.941 | 0.943 | 0.946 |
| RF | 0.963 | 0.961 | 0.965 | 0.965 | 0.961 |
| DT | 0.932 | 0.935 | 0.937 | 0.933 | 0.932 |
| kNN | 0.951 | 0.956 | 0.955 | 0.949 | 0.959 |
| Features: TCP + HTTP | |||||
| SVM | 0.938 | 0.933 | 0.934 | 0.933 | 0.932 |
| RF | 0.952 | 0.953 | 0.956 | 0.955 | 0.951 |
| DT | 0.929 | 0.926 | 0.925 | 0.922 | 0.925 |
| kNN | 0.967 | 0.964 | 0.967 | 0.969 | 0.964 |
@@
Zulkifli et al. [Zulkifli2018] used DT with TCP-based features in their study. They used a small dataset of 500 benign and 200 malicious applications, the latter assembled from two existing malware collections (Drebin and contagiodumpset). Notably, all the applications were executed on a real Android device. They extracted the following features:
@@
To further investigate network traffic features, we reimplemented another study by Wang et al. [Wang2016] involving network data. Wang et al. extracted both TCP and HTTP features from network traffic. Wang et al. used the same set of TCP features as Zulkifli et al., with the exception of average packet size. The features they extracted from HTTP headers were:
@@
Table 26 shows the results, and compares them against those from the original study. For our reimplementation, Table 26 shows the results of the best ML models, which were RFs for both the HTTP and TCP feature sets. We observed no benefit to combining the two feature sets, but, for completeness, we also show results for the other models for both the HTTP and combined feature sets in Table 27. Whilst the results show good discrimination between benign applications and malware, we did not see the exceptionally high level of accuracy reported by Wang et al. for HTTP features, and in fact found these to be less discriminative than TCP features. These conclusions are supported by a positive Kruskal-Wallis test (H = 14.29, p 0.05) and Dunn’s tests.
@@
6.4 Comparison of dynamic modelling approaches
@@
7 Reimplementation and extension of hybrid modelling approaches
| Kandukuru and Sharma Results | Our Results | |||||||||
| DT | 0.956 | - | - | 0.95 | 0.96 | 0.921 | 0.923 | 0.923 | 0.926 | 0.925 |
| RF | - | - | - | - | - | 0.943 | 0.941 | 0.945 | 0.948 | 0.940 |
| SVM | - | - | - | - | - | 0.933 | 0.936 | 0.935 | 0.939 | 0.937 |
| kNN | - | - | - | - | - | 0.936 | 0.938 | 0.939 | 0.941 | 0.937 |
@@ 7 Reimplementation and extension of hybrid modelling approaches
| DT | 0.919 | 0.919 | 0.917 | 0.923 | 0.917 |
| Classifier | Features | Accuracy | Precision | F1-Score | TPR | TNR |
@@
We begin by reimplementing a study by Kandukuru and Sharma [Kandukuru2017] that combined dynamic network traffic features with static permission features. The network traffic features included all seven of the TCP features listed in Section 6.3 and the authors used a malware dataset of 1,200 applications collected from 2010 to 2011 to train a decision tree model. In our reimplementation, we trained all four core ML models, using a dataset of feature vectors comprising 166 permissions and the seven TCP features. Table 28 compares their performance with Kandukuru and Sharma’s results. Since their performance is worse than models trained only on network features (see Tables tables 25 and 26), this suggests that there is no benefit to including information about static permissions.
To further investigate the use of hybrid features, we reimplemented a study based around two frequently used groups of static and dynamic features, permissions and system calls. Kapratwar et al. [Kapratwar2017] carried out experiments using these features both separately and in combination. In the case of system calls, a frequency-based representation was used. For each experiment, the authors trained RF, decision trees, Naïve Bayes, logistic regression, kNN and SVM. However, as with most past hybrid works, Kapratwar et al. used a small dataset of just 103 malware and 97 benign applications. Therefore, a reimplementation of this approach using a larger contemporary dataset is essential to understand the effect of combining two of the most prevalent features in Android malware detection. The dataset used for this reimplementation included 166 permissions and 599 system calls.
@@
8 Best performing models
8.1 Comparison of existing models
Table 31 lists the best classification models seen during this study. Two modelling approaches stand out in terms of the metric scores achieved when discriminating malware from benign applications. TCP network traffic models using random forests () achieved the highest scores in terms of accuracy, precision and TNR. Five-gram opcode models, again using random forests (), achieved the highest scores in terms of F1 and TPR. As noted in Section 4.5, TPR is particularly important for malware detection, since a benign application labelled as malware is less likely to cause as many security problems as malware labelled as benign. It is also notable that 11 of the 12 classifiers listed in Table 31 use random forests, suggesting that this is a good default choice within malware detection. This reflects the wider success of tree-based ensemble models, especially when working with pre-extracted features [grinsztajn2022tree].
@@ 8.1 Comparison of existing models
A notable omission within this table is the presence of any deep learning models. The deep learning approaches that we reimplemented in this study generally performed significantly less well than simpler approaches. At least within the scope of this study, this suggests they are not a good choice of model to use with pre-extracted features, especially given that deep learning models also bring risks of overfitting, responding to spurious patterns within data, and limited transparency [Lones2021], not to mention higher training and inference times — so in general are not likely to be a productive approach for malware detection unless they lead to a significant improvement in performance. This reflects observations made elsewhere about the limitations of deep learning within a feature-based modelling context [grinsztajn2022tree].
| ID | Feature Set | Type | Model | Selection | Accuracy | F1-Score | Precision | TPR | TNR |
| API | Static | RF | MI | 0.967 | 0.969 | 0.967 | 0.963 | 0.970 | |
| API | Static | RF | VT | 0.968 | 0.971 | 0.968 | 0.965 | 0.971 | |
| API | Static | RF | CS | 0.961 | 0.915 | 0.954 | 0.967 | 0.953 | |
| Permissions | Static | RF | MI | 0.930 | 0.932 | 0.929 | 0.931 | 0.931 | |
| Permissions | Static | RF | VT | 0.949 | 0.948 | 0.950 | 0.951 | 0.949 | |
| Opcodes (5-gram) | Static | RF | MI | 0.973 | 0.979 | 0.975 | 0.965 | 0.979 | |
| System Calls (3-gram) | Dynamic | RF | CS | 0.951 | 0.951 | 0.950 | 0.949 | 0.952 | |
| System Calls + API Call Usage | Dynamic | RF | None | 0.961 | 0.964 | 0.963 | 0.957 | 0.965 | |
| TCP | Dynamic | RF | None | 0.977 | 0.969 | 0.981 | 0.978 | 0.971 | |
| TCP + HTTP | Dynamic | kNN | None | 0.967 | 0.964 | 0.967 | 0.969 | 0.964 | |
| TCP + Permissions | Hybrid | RF | None | 0.943 | 0.941 | 0.945 | 0.948 | 0.940 | |
| Permissions + System Calls | Hybrid | RF | MI | 0.948 | 0.948 | 0.949 | 0.935 | 0.961 |
8.2 Comparison of ensemble models